# Optimal upper bounds for anti-canonical volumes of singular toric Fano varieties

### v1

@@ 5. Bounding the dual normalized volumes of Fano polytopes
In this section, we follow the strategy in [5] to derive an upper bound for the dual volumes of minimal Fano polytopes. Fix two positive integers $d\geq 3$ and $q\geq 2$. Let $P$ be a $d$-dimensional minimal $\frac{1}{q}$-lc Fano polytope such that $P$ is not a simplex. Then there exists a decomposition of $P$ into lower dimensional minimal $\frac{1}{q}$-lc Fano simplices as in Theorem 2.7. Keep the same notation as in Theorem 2.7. According to [5, (3.1)], we have the following upper bound for the normalized volume of $P^{*}$:
@@
where for each $1\leq i\leq t$, $\operatorname{Vol}(S_{i}^{*})$ means the normalized volume of $S_{i}^{*}$ in the corresponding $d_{i}$-dimensional $\mathbb{R}$-vector space. By Theorem 4.1, we have $\operatorname{Vol}(S_{i}^{*})\leq B_{d_{i}}$, where $B_{d_{i}}:=\frac{2u_{d_{i},q}^{2}}{q^{d_{i}+1}}$. Our first step is to figure out when the following inequality holds:
$\frac{(d_{1}+\cdots+d_{t})!}{d_{1}!\cdots d_{t}!}\cdot\Pi_{i=1}^{t}B_{d_{i}}\leq B_{d}.$ (5.2)
@@
Fix two positive integers $d\geq 3$ and $q\geq 2$. Let $P$ be a $d$-dimensional minimal $\frac{1}{q}$-lc Fano polytope. Suppose $P$ is not a simplex and $P$ is decomposed into lower dimensional minimal $\frac{1}{q}$-lc Fano simplices as in Theorem 2.7. Keep the same notation as in Theorem 2.7. Suppose the decomposition is not the case when $t=2,d_{1}=d_{2}=d-1$. Then
$\operatorname{Vol}(P^{*})\leq\frac{2u_{d,q}^{2}}{q^{d+1}}.$
@@
Due to (5.1), we only need to prove (5.2) holds. We separate the proof into two parts: $t\geq 3$ and $t=2$.
@@
By (2.2), $d_{i}\leq d-t+1$. $\frac{(d_{1}+\cdots+d_{t})!}{d_{1}!\cdots d_{t}!}\cdot\Pi_{i=1}^{t}B_{d_{i}}$ will increase if $d_{i}$ increase. So to prove (5.2), it is enough to show
$\frac{(t(d-t+1))!}{(d-t+1)!^{t}}\cdot B_{d-t+1}^{t}\leq B_{d}.$ (5.3)
By assumptions $q\geq 2$, $B_{d-t+1}=\frac{2u_{d-t+1,q}^{2}}{q^{d-t+2}}$ and $B_{d}=\frac{2u_{d,q}^{2}}{q^{d+1}}$, thus (5.3) holds if
$\frac{(t(d-t+1))!}{(d-t+1)!^{t}}\cdot\frac{u_{d-t+1,q}^{2t}}{q^{t(d-t+1)-d}}\leq u_{d,q}^{2}$
@@
$\frac{(t(d-t+1))!}{(d-t+1)!^{t}}\cdot\frac{2^{t}\cdot u_{d-t+1,q}^{2t}}{q^{t(d-t+2)}}\leq\frac{2u_{d,q}^{2}}{q^{d+1}}.$
@@
$\frac{(2d-3)!}{(d-1)!(d-2)!}\cdot B_{d-1}\cdot B_{d-2}\leq B_{d}.$
@@
$\frac{(2d-3)!}{(d-1)!(d-2)!}\cdot\frac{2u_{d-2,q}^{2}}{q^{d-2}}\leq(1+u_{d-1,q})^{2}.$ (5.6)
@@
To finish the proof, we need to make a better estimate of this quantity $\frac{(2d-3)!}{(d-1)!(d-2)!}$ when $3\leq d\leq 8$. For $4\leq d\leq 8$, it can be verified by estimating $\frac{(2d-3)!}{(d-1)!(d-2)!}$ appropriately that
$\frac{(2d-3)!}{(d-1)!(d-2)!}\cdot\frac{2u_{d-2,q}^{2}}{q^{d-2}}<K^{2^{d}}<(1+u_{d-1,q})^{2};$
for $d=3$, $\frac{(2d-3)!}{(d-1)!(d-2)!}\cdot\frac{2u_{d-2,q}^{2}}{q^{d-2}}=6q<q^{4}<(1+u_{2,q})^{2}$.
To bound the volume of $P^{*}$ when $P$ is decomposed into two $(d-1)$-dimensional minimal $\frac{1}{q}$-lc Fano simplices, we apply the result figured out through integration in [5].
Lemma 5.3 (cf. [5, $(6.1),(6.2)$]).
Fix two integers $d\geq 3$ and $q\geq 2$. Let $P$ be a $d$-dimensional minimal $\frac{1}{q}$-lc Fano polytope. Suppose $P$ is not a simplex and is decomposed into two $d-1$-dimensional minimal $\frac{1}{q}$-lc Fano simplices. Keep the same notation as in Theorem 2.7. Suppose the barycentric coordinates of ${\mathbf{0}}$ with respect to $S_{i}$ are $\beta_{i,1},\cdots,\beta_{i,d}$ for each $i=1,2$. Then
@@
Fix two integers $d\geq 3$ and $q\geq 2$. Let $P$ be a $d$-dimensional minimal $\frac{1}{q}$-lc Fano polytope. Suppose $P$ is not a simplex and is decomposed into two $(d-1)$-dimensional minimal $\frac{1}{q}$-lc Fano simplices. Then
$\operatorname{Vol}(P^{*})\leq\frac{2u_{d,q}^{2}}{q^{d+1}}.$
@@
$\frac{1}{\beta_{i,1}\cdots\beta_{i,d}}\leq\frac{u_{d,q}^{2}}{q^{d+1}}.$
So we get the conclusion. ∎
@@ 6. Proof of main results
Let $P$ be a $d$-dimensional $\frac{1}{q}$-lc Fano polytope. Then $P$ contains a $d$-dimensional minimal $\frac{1}{q}$-lc Fano polytope $P^{\prime}$. Thus by Propositions 5.2 and 5.4, we have $\operatorname{Vol}(P^{*})\leq\operatorname{Vol}(P^{\prime*})\leq\frac{2u_{d,q}^{2}}{q^{d+1}}$. The remaining conclusion follows from Example 4.3 and Proposition 4.4. ∎
@@
This follows from Theorem 2.9, Theorem 1.2 and Example 4.3. ∎
@@
By assumption, $\frac{1}{q}S\cap{\mathbf{N}}=\{{\mathbf{0}}\}$. By abuse of notation, we also assume $\beta_{1}\geq\cdots\geq\beta_{d+1}$ are the barycentric coordinates of ${\mathbf{0}}$ with respect to $S$. Then by applying Theorem 2.3, we have
$\operatorname{Vol}(\frac{1}{q}S)\leq\frac{1}{\beta_{1}\cdots\beta_{d}}.$
@@
$\operatorname{Vol}(S)$ ${}\leq\frac{q^{d}}{\beta_{1}\cdots\beta_{d}}$
${}\leq\frac{2u_{d,q}^{2}}{q}.$
To conclude the proof, we only need to show that $\operatorname{int}(\frac{1}{q}S)\cap{\mathbf{N}}=\{{\mathbf{0}}\}$, where $S=\operatorname{conv}((1+u_{1,q})e_{1}-qe,\cdots,(1+u_{d-1,q})e_{d-1}-qe,-u_{d,q}e_{d}-qe,u_{d,q}e_{d}-qe)$. Moreover, we claim that $\frac{1}{q}=\max\{t|\operatorname{int}(tS)\cap{\mathbf{N}}=\{{\mathbf{0}}\}\}$.
Take a real number $q^{\prime}$ such that $0<q^{\prime}\leq q$. Suppose $v\in\operatorname{int}(\frac{1}{q^{\prime}}S)\cap{\mathbf{N}}$ has barycentric coordinates $\lambda_{1},\cdots,\lambda_{d+1}$. Then for $i\in\{1,\cdots,d-1\}$, $\lambda_{i}=\frac{q^{\prime}z_{i}+q}{1+u_{i,q}}$ for some $z_{i}\in{\mathbb{Z}}$, and $\lambda_{d+1}-\lambda_{d}=\frac{q^{\prime}z}{u_{d,q}}$ for some $z\in{\mathbb{Z}}$.
If $q^{\prime}=q$, then $\lambda_{i}>0$ implies $\lambda_{i}\geq\frac{q}{1+u_{i,q}}$ for $i\in\{1,\cdots,d-1\}$. By (1.1) and the assumption that $v$ is an interior lattice point of $\frac{1}{q}S$, we must have $\lambda_{i}=\frac{q}{1+u_{i,q}}$ for $i\in\{1,\cdots,d-1\}$. And then $\lambda_{d}+\lambda_{d+1}=\frac{q}{u_{d,q}}$, which implies $\lambda_{d}=\lambda_{d+1}=\frac{q}{2u_{d,q}}$. So we conclude that $\operatorname{int}(\frac{1}{q}S)\cap{\mathbf{N}}=\{{\mathbf{0}}\}$.
If $0<q^{\prime}<q$, then we can show that $e_{d}\in\operatorname{int}(\frac{1}{q^{\prime}}S)\cap{\mathbf{N}}$ by taking $\lambda_{i}=\frac{q}{1+u_{i,q}}$ for $i\in\{1,\cdots,d-1\}$, $\lambda_{d}=\frac{q-q^{\prime}}{2u_{d,q}}$ and $\lambda_{d+1}=\frac{q+q^{\prime}}{2u_{d,q}}$.

### v2

@@ 5. Bounding the dual normalized volumes of Fano polytopes
In this section, we follow the strategy in [6] to derive an upper bound for the dual normalized volumes of minimal Fano polytopes. Fix two positive integers $d\geq 3$ and $q\geq 2$. Let $P$ be a $d$-dimensional minimal $\frac{1}{q}$-lc Fano polytope such that $P$ is not a simplex. Then there exists a decomposition of $P$ into lower dimensional $\frac{1}{q}$-lc Fano simplices as in Theorem 2.7. Keep the same notation as in Theorem 2.7. According to the monotonicity of the normalised volume (cf. [6, (3.1)]), we have the following upper bound for the normalized volume of $P^{*}$:
@@
where for each $1\leq i\leq t$, $\operatorname{Vol}(S_{i}^{*})$ means the normalized volume of $S_{i}^{*}$ in the corresponding $d_{i}$-dimensional $\mathbb{R}$-vector space. By Theorem 4.1, we have $\operatorname{Vol}(S_{i}^{*})\leq\frac{2u_{d_{i},q}^{2}}{q^{d_{i}+1}}$. Our first step is to figure out when the following inequality holds:
$\frac{(d_{1}+\cdots+d_{t})!}{d_{1}!\cdots d_{t}!}\cdot\Pi_{i=1}^{t}\frac{2u_{d_{i},q}^{2}}{q^{d_{i}+1}}<\frac{2u_{d,q}^{2}}{q^{d+1}}.$ (5.2)
@@
Fix two positive integers $d\geq 3$ and $q\geq 2$. Let $P$ be a $d$-dimensional minimal $\frac{1}{q}$-lc Fano polytope. Suppose $P$ is not a simplex and $P$ is decomposed into lower dimensional $\frac{1}{q}$-lc Fano simplices as in Theorem 2.7. Keep the same notation as in Theorem 2.7. Suppose the decomposition is not the case when $t=2,d_{1}=d_{2}=d-1$. Then
$\operatorname{Vol}(P^{*})<\frac{2u_{d,q}^{2}}{q^{d+1}}.$
@@
Due to (5.1) and Theorem 4.1, we only need to prove (5.2) holds. We separate the proof into two parts: $t\geq 3$ and $t=2$.
@@
By (2.2), $d_{i}\leq d-t+1$. $\frac{(d_{1}+\cdots+d_{t})!}{d_{1}!\cdots d_{t}!}\cdot\Pi_{i=1}^{t}\frac{2u_{d_{i},q}^{2}}{q^{d_{i}+1}}$ will increase if $d_{i}$ increase. So to prove (5.2), it is enough to show
$\frac{(t(d-t+1))!}{(d-t+1)!^{t}}\cdot(\frac{2u_{d-t+1,q}^{2}}{q^{d-t+2}})^{t}<\frac{2u_{d,q}^{2}}{q^{d+1}}.$ (5.3)
Since $q\geq 2$, (5.3) holds if
$\frac{(t(d-t+1))!}{(d-t+1)!^{t}}\cdot\frac{u_{d-t+1,q}^{2t}}{q^{t(d-t+1)-d}}<u_{d,q}^{2}$
@@
$\frac{(t(d-t+1))!}{(d-t+1)!^{t}}\cdot\frac{2^{t}\cdot u_{d-t+1,q}^{2t}}{q^{t(d-t+2)}}<\frac{2u_{d,q}^{2}}{q^{d+1}}.$
@@
$\frac{(2d-3)!}{(d-1)!(d-2)!}\cdot\frac{2u_{d-1,q}^{2}}{q^{d}}\cdot\frac{2u_{d-2,q}^{2}}{q^{d-1}}<\frac{2u_{d,q}^{2}}{q^{d+1}}.$
@@
$\frac{(2d-3)!}{(d-1)!(d-2)!}\cdot\frac{2u_{d-2,q}^{2}}{q^{d-2}}<(1+u_{d-1,q})^{2}.$ (5.6)
@@
To finish the proof, we need to make a better estimate of this quantity $\frac{(2d-3)!}{(d-1)!(d-2)!}$ when $3\leq d\leq 8$. For $3\leq d\leq 8$, it can be verified by estimating $\frac{(2d-3)!}{(d-1)!(d-2)!}$ appropriately that
$\frac{(2d-3)!}{(d-1)!(d-2)!}\cdot\frac{2u_{d-2,q}^{2}}{q^{d-2}}<K^{2^{d}}<(1+u_{d-1,q})^{2}.$
To bound the volume of $P^{*}$ when $P$ is decomposed into two $(d-1)$-dimensional $\frac{1}{q}$-lc Fano simplices, we apply the result figured out through integration in [6].
Lemma 5.3 (cf. [6, $(6.1),(6.2)$]).
Fix two integers $d\geq 3$ and $q\geq 2$. Let $P$ be a $d$-dimensional minimal $\frac{1}{q}$-lc Fano polytope. Suppose $P$ is not a simplex and is decomposed into two $(d-1)$-dimensional $\frac{1}{q}$-lc Fano simplices. Keep the same notation as in Theorem 2.7. Suppose the barycentric coordinates of ${\mathbf{0}}$ with respect to $S_{i}$ are $\beta_{i,1},\cdots,\beta_{i,d}$ for each $i=1,2$. Then
@@
Furthermore, if the equality holds, then for each $i=1,2$, there exists two of $\beta_{i,1},\cdots,\beta_{i,d}$ are equal.
@@
Fix two integers $d\geq 3$ and $q\geq 2$. Let $P$ be a $d$-dimensional minimal $\frac{1}{q}$-lc Fano polytope. Suppose $P$ is not a simplex and is decomposed into two $(d-1)$-dimensional $\frac{1}{q}$-lc Fano simplices. Then
$\operatorname{Vol}(P^{*})<\frac{2u_{d,q}^{2}}{q^{d+1}}.$
@@
$\operatorname{Vol}(P^{*})\leq\frac{1}{\beta_{1,1}\cdots\beta_{1,d}}+\frac{1}{\beta_{2,1}\cdots\beta_{2,d}}.$ (5.7)
@@
$\frac{1}{\beta_{i,1}\cdots\beta_{i,d}}\leq\frac{u_{d,q}^{2}}{q^{d+1}}.$ (5.8)
So we get the conclusion by noting that the equality in (5.7) and the equality in (5.8) cannot hold simultaneously. ∎
@@ 6. Proof of main results
If $P_{X}$ is not a simplex, then $(-K_{X})^{d}=\operatorname{Vol}(P_{X}^{*})<\frac{2u_{d,q}^{2}}{q^{d+1}}$ by Theorem 2.9, Proposition 5.2 and Proposition 5.4. If $P_{X}$ is a simplex, then the conclusion follows from Theorem 4.2. ∎
@@
We only need to consider the case when $P$ is a $d$-dimensional minimal $\frac{1}{q}$-lc Fano polytope. Then the conclusions follow from Theorem 4.1, Proposition 5.2 and Proposition 5.4. ∎
@@
By assumption, $\frac{1}{q}S\cap{\mathbb{Z}}^{d}=\{{\mathbf{0}}\}$. By abuse of notation, we also assume $\beta_{1}\geq\cdots\geq\beta_{d+1}$ are the barycentric coordinates of ${\mathbf{0}}$ with respect to $S$. Then by applying Theorem 2.3, we have
$\operatorname{vol}_{{\mathbb{Z}}}(\frac{1}{q}S)\leq\frac{1}{d!\beta_{1}\cdots\beta_{d}}.$
@@
$\operatorname{vol}_{{\mathbb{Z}}}(S)$ ${}\leq\frac{q^{d}}{d!\beta_{1}\cdots\beta_{d}}$
${}\leq\frac{2u_{d,q}^{2}}{d!q}.$
To conclude the proof, we only need to show that $\operatorname{int}(\frac{1}{q}S)\cap{\mathbb{Z}}^{d}=\{{\mathbf{0}}\}$, where $S=\operatorname{conv}((1+u_{1,q})e_{1}-qe,\cdots,(1+u_{d-1,q})e_{d-1}-qe,-u_{d,q}e_{d}-qe,u_{d,q}e_{d}-qe)$. Moreover, we claim that $\frac{1}{q}=\max\{t|\operatorname{int}(tS)\cap{\mathbb{Z}}^{d}=\{{\mathbf{0}}\}\}$.
Take a real number $q^{\prime}$ such that $0<q^{\prime}\leq q$. Suppose $v\in\operatorname{int}(\frac{1}{q^{\prime}}S)\cap{\mathbb{Z}}^{d}$ has barycentric coordinates $\lambda_{1},\cdots,\lambda_{d+1}$, i.e.,
$\Sigma_{i=1}^{d+1}\lambda_{i}=1,\lambda_{i}>0,\forall i\in\{1,\cdots,d+1\},$
and
$v=\Sigma_{i=1}^{d-1}\frac{\lambda_{i}}{q^{\prime}}((1+u_{i,q})e_{i}-qe)+\frac{\lambda_{d}}{q^{\prime}}(-u_{d,q}e_{d}-qe)+\frac{\lambda_{d+1}}{q^{\prime}}(u_{d,q}e_{d}-qe)\in{\mathbb{Z}}^{d}.$
Then for $i\in\{1,\cdots,d-1\}$, $\lambda_{i}=\frac{q^{\prime}z_{i}+q}{1+u_{i,q}}$ for some $z_{i}\in{\mathbb{Z}}$, and $\lambda_{d+1}-\lambda_{d}=\frac{q^{\prime}z}{u_{d,q}}$ for some $z\in{\mathbb{Z}}$.
If $q^{\prime}=q$, then $\lambda_{i}>0$ implies $\lambda_{i}\geq\frac{q}{1+u_{i,q}}$ for $i\in\{1,\cdots,d-1\}$. By (1.1) and the assumption that $v$ is an interior lattice point of $\frac{1}{q}S$, we must have $\lambda_{i}=\frac{q}{1+u_{i,q}}$ for $i\in\{1,\cdots,d-1\}$. And then $\lambda_{d}+\lambda_{d+1}=\frac{q}{u_{d,q}}$, which implies $\lambda_{d}=\lambda_{d+1}=\frac{q}{2u_{d,q}}$. So we conclude that $\operatorname{int}(\frac{1}{q}S)\cap{\mathbb{Z}}^{d}=\{{\mathbf{0}}\}$.
If $0<q^{\prime}<q$, then we can show that $e_{d}\in\operatorname{int}(\frac{1}{q^{\prime}}S)\cap{\mathbb{Z}}^{d}$ by taking $\lambda_{i}=\frac{q}{1+u_{i,q}}$ for $i\in\{1,\cdots,d-1\}$, $\lambda_{d}=\frac{q-q^{\prime}}{2u_{d,q}}$ and $\lambda_{d+1}=\frac{q+q^{\prime}}{2u_{d,q}}$.